3) Assuming that firms compete a la Cournot, that all firms have the same marginal cost, and

that demand is linear, when is price most sensitive to changes in marginal cost: in a market

with very few firms or in a market with many firms? Show this formally. [Hint: assume

demand 𝑝 = π‘Ž – 𝑏𝑄]

b) Consider a Bertrand duopoly with differentiated products. Demand curves are given by

𝑝1 = 600 βˆ’ 2π‘ž1 βˆ’ π‘ž2

𝑝2 = 600 βˆ’ π‘ž1 βˆ’ 2π‘ž2

Suppose that the cost functions are given by (π‘ži) = 60π‘ži , for 𝑖 = 1, 2. Find the equilibrium

outputs, the prices and the profits.


1
Expert's answer
2021-10-20T10:05:43-0400

Solution:

a.). The price is most sensitive to changes in marginal cost when the number of firms increase.

If the oligopoly is symmetric, meaning that all firms have identical products and cost conditions, the degree to which price exceeds marginal cost is inversely related to the number of firms.

Β 

b.). TR1 = (600 – 2q1 – q2) q1 = 600q1 – 2q12 – q2q1

Profit1 = (600q1 – 2q12 – q2q1) – 60q1

"\frac{\partial \pi 1} {\partial q1}" = 600 – 4q1 – q2 – 60

600 – 4q1 – q2 – 60 = 0

4q1 = 540 – q2

q1 = 135 – 0.25q2


TR2 = (600 – q1 – 2q2) q2 = 600q2 – q1q2 – 2q22

Profit1 = (600q2 – q1q2 – 2q22) – 60q2

"\frac{\partial \pi 2} {\partial q2}" = 600 – q1 – 4q2 – 60

600 – q1 – 4q2 – 60 = 0

4q2 = 540 – q1

q2 = 135 – 0.25q1

Β 

q1 = 135 – 0.25q2

q1 = 135 – 0.25(135 – 0.25q1)

q1 = 135 – 33.75 + 0.0625q1

q1 = 101.25 + 0.0625q1

q1 – 0.025q1 = 101.25

0.9375q1 = 101.25

q1 = 108

q2 = 108

Equilibrium quantity = 108


P1 = 600 – 2q1 – q2 = 600 – 2(108) – 108 = 276

Equilibrium price = 276

Β 

Profit = TR – TC

Profit = (P "\times" Q) – (60q1)

Profit = (276"\times"108) – (60 "\times"108)

Profit = 29,808 – 6,480 = 23,328

Profit = 23,328